Session: Oct 2008 (session 3)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Bo Xia at Fri Oct 17 19:07:09 2008) Success :

Urumqi (Hua Zhang at Sat Oct 18 04:32:22 2008) Success :

Onsala (Michael Lindqvist at Sat Oct 18 18:30:25 2008) Success :
No known problems.

Effelsberg (Uwe Bach at Mon Oct 20 07:33:11 2008) Minor failures : LATE
FS started with wrong time. Missed the first 20 minutes.

Yebes (Pablo de Vicente at Mon Oct 20 09:14:45 2008) Success :

Westerbork (Geert Kuper at Wed Oct 22 21:00:37 2008) Minor failures : PHAS
Wb used in phased array mode. Observation was done using TADUMax and Mark5B X/Y phase difference not optimal Especially 8419 USB and 8427 LSB Disk2file command only worked manually using byteoffsets. The resulting fringefindfile had to be repaired by the Jive staff.

Medicina (Andrea Orlati at Thu Oct 23 07:27:22 2008) Minor failures : RECE
Problems with noise diode. Experiment started with no valid tsys information. Forced to stop in order to solve the problem. Lost scans from 13:16 to 13:32 UT

Noto (P. Cassaro at Fri Oct 31 10:34:54 2008) Success :
No known problems